### v2

@@ Appendix A Results Across all 4 LLMs
Figure 13: Temperature hyperparameter does not solve flatness for Wizard Vicuna Uncensored. Same interpretation as Fig. 5: comparison of human in-group diversity to Wizard Vicuna Uncensored generations varying levels of temperature settings, where by 1.8 the responses become incoherent. At this setting even though the unique n-gram metric shows the LLM surpassing humans in diversity, this is only due to the incoherence as under no other semantic metric is human diversity reached.
@@
Figure 16: Average multiple choice responses for LLMs and human participants. LLM responses are indicated in colors, human in-group in black, and human out-group in gray. For R1-Contingent, 1 to 5 represents how challenging it is to have that demographic identity; R2-Relevant 1 to 5 represents conservative to liberal political opinion; R3-Subj represents level of toxicity detected for race and level of positive sentiment detected for gender. Along the demographic axes of race and intersectional, out-group human participants tend to overinflate the difficulty of being in that group compared to in-group human participants, and LLMs even further inflate the difficulty beyond that of the out-group. GPT-4 inflates the difficulty more for White men and White women compared to Black men and Black women.

### v3

@@ Appendix A Results Across all 4 LLMs
Figure 13: Temperature hyperparameter does not solve flatness for Wizard Vicuna Uncensored. Same interpretation as Fig. 4 in the main text: comparison of human in-group diversity to Wizard Vicuna Uncensored generations varying levels of temperature settings, where by 1.8 the responses become incoherent. There are 100 samples or generations per scenario, and 95% confidence intervals are generated via bootstrapping. At this setting even though the unique n-gram metric shows the LLM surpassing humans in diversity, this is only due to the incoherence as under no other semantic metric is human diversity reached.
@@
Figure 16: Average multiple choice responses for LLMs and human participants. LLM responses are indicated in colors, human in-group in black, and human out-group in gray. For R1-Contingent, 1 to 5 represents how challenging it is to have that demographic identity; R2-Relevant 1 to 5 represents conservative to liberal political opinion; R3-Subj represents level of toxicity detected for race and level of positive sentiment detected for gender. There are 100 generations or samples per scenario, and 95% confidence intervals are generated through bootstrapping. Along the demographic axes of race and intersectional, out-group human participants tend to overinflate the difficulty of being in that group compared to in-group human participants, and LLMs even further inflate the difficulty beyond that of the out-group. GPT-4 inflates the difficulty more for White men and White women compared to Black men and Black women.